# Axion Icebergs: Clockwork ALPs at hadron colliders

### v1

@@ 4.2 Intermediate mass ALPs
With an intermediate mass range set by the parameters $m=35$ GeV and $f=1000$ GeV, the analysis particulars in this case are largely the same as described for benchmark I. With the ALP masses in this case spanning the range $35-105$ GeV, a sizable fraction of the spectrum still falls in the low mass region for which we require the final state photons to be boosted due to reasons delineated for benchmark I. We, therefore, choose to adopt the same strategy as in benchmark I for the collider simulation and consider a signal consisting of CW axions produced in association with upto two jets (and the axions further decaying to two photons). Photons and jets are reconstructed using the same criteria as in the case of Benchmark point I (see Table 1).
4.2.1 Benchmark Point II(a)
@@
The $p_{T}$ distributions for the leading and the sub-leading photons thus obtained for the signal events alongwith the corresponding distributions for the jet $p_{T}$’s are shown in Fig.9 (left), assuming a bin size of 1 GeV. The distributions are similar in nature to what was obtained for benchmark point I, albeit with a shift in the photon distributions towards higher $p_{T}$ values due to the ALPs being heavier in this case. The attendant diphoton invariant mass distribution is shown in Fig.9 (right). As for the SM backgrounds, for the mass range $m_{\gamma\gamma}\in[10,80]$ GeV, once again we adopt the ATLAS analysis ATLAS:2022abz . For the $[80,110]$ GeV window, in the absence of an appropriate analysis [^16] The ATLAS search for diphoton resonances in the mass range $m_{\gamma\gamma}=66-110$ GeV ATLAS:2024bjr does provide a background estimation in the region of interest. However, the event selection criteria employed therein is different than that in the low mass analysis that we refer to, which makes it difficult to match the background events corresponding to the low and intermediate mass ranges in a meaningful way., we choose to overestimate the background, by holding it at the level at $m_{\gamma\gamma}=80$ GeV. Thus, our significance estimate would be somewhat conservative.
@@
Figure 9: BP-II(a): event distributions, after applying acceptance cuts alone, vide. Table 1. (left) $p_{T}$ distributions for the leading photon ($\gamma_{1}$), sub-leading photon ($\gamma_{2}$), leading jet ($j_{1}$) and the sub-leading jet ($j_{2}$) and (right) the two-photon invariant mass.
@@
To obtain the signal diphoton distribution in the fiducial region we employ the selection cuts mentioned in Table 2. As for benchmark I, the kinematic cuts substantially reduce the number of events in the signal distribution, shown in Fig.10. Given the characteristic mass-splittings of this benchmark, we see that a large portion of the spectrum (consisting of 28 ALPs) in the diphoton distribution could potentially be resolved at the detector. Now, a bin-wise significance estimation, under the conservative assumption of the background events beyond $m_{\gamma\gamma}\approx 80$ GeV, leads to a distribution as shown in Fig.11 with a cumulative significance $S_{c,138}=2.12\sigma$ over the signal region.
4.2.2 Benchmark Point II(b)
@@
4.3 Heavy ALPs
For ALPs that are significantly heavier than considered hitherto, the decay photons would be expected to carry sufficient energy (and $p_{T}$) for the event to be triggered even without any additional jet. An example is afforded by the aforementioned Benchmark III, wherein the ALP masses span the range $\sim 150-450$ GeV. However, although extra jets are not needed for triggering, events with such jets do contribiute significantly to the signal cross sections. Hence, we define the signal as a semi-inclusive one composed of a pair of energetic photons with upto two additional jets [^17] The inclusion of three or more jets does not improve the signal to noise ratio.. The corresponding $K$-factors, for this mass range, are taken from refs.Ahmed:2016otz ; Boughezal:2015dra .
@@ 4.3 Heavy ALPs
Table 3: Acceptance cuts for the final state objects in benchmark III ATLAS:2021uiz .
The background, thus, would receive contributions from essentially the same channels as assumed in the low mass case. The signal profile being different, we, though, would need to adequately tune the assessment, and, in this, we are guided by the ATLAS diphoton analysis ATLAS:2021uiz germane to this mass range. Post showering and hadronization, using Pythia8, we take for the identification and isolation of photons in the Delphes detector simulation a cone size of $\Delta R=0.4$ around the photon candidate. The requirement on the parameter $p_{T}^{\rm ratio}$, as defined earlier, is now reset to $p_{T}^{\rm ratio}<0.12$. The complete photon and jet identification criteria, as adapted from the ATLAS analysis ref.ATLAS:2021uiz , are displayed in Table 3. The $p_{T}$ distributions of the photons and the jets for the generated signal [^18] In contrast with benchmarks I and II, events without jets in the final state also contribute significantly in this case and, therefore, $N_{(}2j)<N_{(}1j)<N_{Tot}$. events are shown in Fig.15, adopting a bin size of $16$ GeV as in the ATLAS searchATLAS:2021uiz .
Fig.15 (right) shows the corresponding diphoton invariant mass distribution. Taking cue from ref.ATLAS:2021uiz , we then apply the selection cuts mentioned in Table 4 in order to define the fiducial volume in the phase space.
@@
$pp\rightarrow a_{n}\,$ $|\eta_{\gamma}|<2.37$ (excluding barrel-to-endcap region $1.521.37<|\eta_{\gamma}|<$ 1.52),
@@
Table 4: Event selection cuts applied to form the fiducial signal regions for benchmark III ATLAS:2021uiz .
The resulting diphoton invariant mass distribution is shown in Fig.16, whereas Fig.17 shows the corresponding bin-wise significance (with the background profile adopted from ref.ATLAS:2021uiz ) over the diphoton signal region with a cumulative significance $S_{c,138}\approx 1.58\sigma$. For this case, the projected enhancement for a $300\,\rm fb^{-1}$ luminosity reach is $S_{c,300}\approx 2.33\sigma$. Thus, in contrast to benchmarks I and II, the high mass scenario will not be readily accessible in the LHC’s Run 3. It is perhaps worth speculating, though, that benchmark III, with heavy ALPs, could potentially reach the discovery threshold during the forthcoming high luminosity phase of the LHC with a projected luminosity enhancement by a factor of nearly 20 times the current value.
Figure 15: BP-III: event distributions, after applying acceptance cuts alone, vide. Table 3. (left) $p_{T}$ distributions for the leading photon ($\gamma_{1}$), sub-leading photon ($\gamma_{2}$), leading jet ($j_{1}$) and the sub-leading jet ($j_{2}$) and (right) the two-photon invariant mass.
@@
5 Vector-like quarks and heavy scalars
Beyond the low energy spectrum of the pseudoscalars, the model, as described in section 2, also contains a heavy quark with the $SU(3)_{c}\times SU(2)_{L}\times U(1)_{Y}$ charge assignment $(3,1,2/3)$ as well as $N+1$ heavy radial scalar singlets (the partners of the pseudoscalars). Although not germane to the main objectives of our analysis, it is worth outlining here the dynamics of both the VLQ and the heavy scalars so as to establish the consistency of the model, especially in view of the fact that the assumed mass scales (characterised by the SSB scale $f$) of these heavy particles are, in principle, accessible at the LHC.
Considering that $m^{2}\ll f^{2}$ in all the benchmarks, the off-diagonal quadratic terms for the radial scalars in the Lagrangian (eq.(4)) are hierarchically smaller than the diagonal mass terms governed by the coupling $\lambda$, which is presumed to assume $\mathcal{O}(1)$ values. Consequently, all the radial scalars are nearly degenerate with masses $\sim\sqrt{3\lambda}f$. To simplify matters, we assume the other new quartic coupling, viz. $\lambda_{\Phi H}$, is also small, thereby automatically relaxing constraints from the stability of the SM Higgs potential, triviality, etc. Similar to the pseudoscalars, these heavy scalars have effective interactions with the gluons, photons and the EW vector bosons with the respective couplings given by $g_{\pi VV}/\xi$, where $g_{\pi VV}$ represent the pseudoscalar couplings as listed in eq.(16).
@@ 5 Vector-like quarks and heavy scalars
$\left(\Psi_{L}\quad u^{(3)}_{L}\right)^{T}=U\,\left(T_{L}\quad t_{L}\right)^{T}\ ,\qquad{\rm and}\qquad\left(\Psi_{R}\quad u^{(3)}_{R}\right)^{T}=V\,\left(T_{R}\quad t_{R}\right)^{T}\ ,$ (22)
@@
$\begin{array}[]{rcl}{\cal L}_{T-V}&=&\displaystyle\left(\frac{g}{\sqrt{2}}U_{tT}\right)W_{\mu}\,\bar{b}_{L}\gamma^{\mu}T_{L}\;+\left(\frac{g}{2c_{w}}U^{*}_{tt}U_{tT}\right)Z_{\mu}\,\bar{t}_{L}\gamma^{\mu}T_{L}\\[8.61108pt] &+&\displaystyle\frac{g}{c_{w}}\,Z_{\mu}\,\bar{T}\gamma^{\mu}\left[-\,\frac{2}{3}s_{w}^{2}+\frac{g}{2c_{w}}U^{*}_{Tt}U_{tT}P_{L}\right]T+\mbox{h.c.}\end{array}$ (23)
@@
$\begin{split}\mathcal{L}_{T-\phi}&=-\frac{1}{\sqrt{2}}\bar{t}_{L}\left[\left(\lambda_{\Psi}U^{*}_{tT}V_{TT}+y^{\prime}_{\Psi}U^{*}_{tT}V_{tT}\right)\phi_{N}+\left(\lambda_{h}U^{*}_{tt}V_{tT}+y_{\Psi}U^{*}_{tt}V_{TT}\right)h\right]T_{R}\\ &-\frac{1}{\sqrt{2}}\bar{T}_{L}\left[\left(\lambda_{\Psi}U^{*}_{TT}V_{Tt}+y^{\prime}_{\Psi}U^{*}_{TT}V_{tt}\right)\phi_{N}+\left(\lambda_{h}U^{*}_{Tt}V_{tt}+y_{\Psi}U^{*}_{Tt}V_{Tt}\right)h\right]t_{R}+\mbox{h.c.}.\end{split}$ (24)
To obtain the pseudoscalar FCNCs, it is perhaps the most convenient to make use of the shift symmetry and rescale the field $\Psi_{L}\to e^{i\xi\pi_{N}/f}\Psi_{L}$ in eq.(10). This, in turn, results in a pseudovector FCNC term in the full Lagrangian through the kinetic term of the field $\Psi_{L}$, namely,
@@
with $\pi_{N}$ being expressed in terms of the mass eigenstates $a_{n}$ through $\pi_{N}=\sum C^{-1}_{Nk}a_{k}$ with the matrix $C$ as defined in eq.(6). For $f\gtrsim 1000$ GeV and $m_{T}\sim\lambda_{\Psi}f/\sqrt{2}\,>(m_{W,Z,h,\phi}+m_{t})$, the primary decay channels for the VLQ are clearly $T\to b\,W$, $T\to t\,Z$, $T\to t\,h$, $T\to t\,\phi_{N}$ (if kinematically allowed) and $T\to t\,a_{n}$ (see appendix B for the decay width expressions Atre:2011ae ). Note that the decays to the new scalars and pseudoscalars appear in addition to the conventional VLQ decay modes to the SM particles and, therefore, it is worth comparing the branching fractions pertaining to the different channels. As this needs specifying a few further parameters, we begin by doing so for each of the benchmark points:
BP-I and II: We assume $\lambda_{\Psi}=2.2$, $y_{\Psi}=\epsilon y^{\prime}_{\Psi}$, $\lambda=1.8$ and $y^{\prime}_{\Psi}\lesssim 0.1$ with $\epsilon=0.1$. This choice of $\lambda_{\Psi}$ is motivated by the experimental lower bound on the VLQ mass ($m_{T}\gtrsim 1500$ GeV) to be discussed below. However, as can be expected from such a large value of the coupling, the evolution is fast and an examination of the three-loop renormalization group equations Buttazzo:2013uya leads to the conclusion that the theory becomes a very strongly coupled one at a scale $\mu\sim 10$ TeV, signalling that new physics must take over well before this scale. Given that the clockwork model is not a UV-complete one, this, per se, may still be overlooked. However, we return to this point in the next section.
While in eqs.(23 & 24), we have not listed any alterations in the SM couplings of the top-quark, it is obvious that certain changes would be wrought. However, these changes are only higher-order in the $T$$u_{3}$ mixing $U_{tT}$, and with the latter not being large, are well below the current sensitivity limits (the strongest being that for the SM CKM element $V_{tb}$ ParticleDataGroup:2024cfk ), whether from flavour physics or from top-decay.
The constraints on the VLQ sector would, thus, come from direct observations at colliders. At the large hadron collider, the overwhelmingly leading production mechanism is the QCD-driven one [^19] With $U_{tT}$ not being large, single production is suppressed and the consequent bounds ATLAS:2023pja are relaxed.. Once pair-produced, the $T$s would decay promptly. Table 5 lists the leading VLQ branching ratios for the three benchmark points [^20] For the chosen set of couplings, a decay of $T$ to radial scalars is kinematically forbidden. For any phenomenologically viable set of couplings, the branching fractions remain much smaller than those listed in the table. It can be readily ascertained that the branching fractions have a very small dependence on the free parameter $y^{\prime}_{\Psi}$.
@@
$m_{T}$ lower limit 1540 GeV CMS:2022fck ; Alves:2023ufm 1500 GeV $\approx$ 1540 GeV
Table 5: VLQ branching fractions for BP-I,II and III (The SM column refers to the assumed branching fractions for VLQ searches ATLAS:2023pja ; CMS:2022fck ). The bottom row displays the resulting lower bounds on $m_{T}$, as derived from the limits on the VLQ pair-production cross-section presented in ref.CMS:2022fck

### v2

@@ IV.2 Intermediate mass ALPs
With an intermediate mass range set by the parameters $m=35$ GeV and $f=1000$ GeV, the analysis particulars in this case are largely the same as described for benchmark I. With the ALP masses in this case spanning the range 3510535-105 GeV, a sizable fraction of the spectrum still falls in the low mass region for which we require the final state photons to be boosted due to reasons delineated for benchmark I. We, therefore, choose to adopt the same strategy as in benchmark I for the collider simulation and consider a signal consisting of CW axions produced in association with upto two jets (and the axions further decaying to two photons). Photons and jets are reconstructed using the same criteria as in the case of Benchmark point I (see Table 1).
IV.2.1 Benchmark Point II(a)
@@
The $p_{T}$ distributions for the leading and the sub-leading photons thus obtained for the signal events alongwith the corresponding distributions for the jet $p_{T}$’s are shown in Fig.9 (top), assuming a bin size of 1 GeV. The distributions are similar in nature to what was obtained for benchmark point I, albeit with a shift in the photon distributions towards higher $p_{T}$ values due to the ALPs being heavier in this case. The attendant diphoton invariant mass distribution is shown in Fig.9 (bottom). As for the SM backgrounds, for the mass range $m_{\gamma\gamma}\in[10,80]$ GeV, once again we adopt the ATLAS analysis Aad et al. (2023a). For the $[80,110]$ GeV window, in the absence of an appropriate analysis [^26] The ATLAS search for diphoton resonances in the mass range $m_{\gamma\gamma}=66-110$ GeV Aad et al. (2024b) does provide a background estimation in the region of interest. However, the event selection criteria employed therein is different than that in the low mass analysis that we refer to, which makes it difficult to match the background events corresponding to the low and intermediate mass ranges in a meaningful way., we choose to overestimate the background, by holding it at the level of $m_{\gamma\gamma}=80$ GeV. Thus, our significance estimate would be somewhat conservative.
@@
Figure 9: BP-II(a): event distributions, after applying acceptance cuts alone, vide. Table 1. (top) $p_{T}$ distributions for the leading photon ($\gamma_{1}$), sub-leading photon ($\gamma_{2}$), leading jet ($j_{1}$) and the sub-leading jet ($j_{2}$) and (bottom) the two-photon invariant mass.
@@
To obtain the signal diphoton distribution in the fiducial region, we employ the selection cuts mentioned in Table 2. As for benchmark I, the kinematic cuts substantially reduce the number of events in the signal distribution, shown in Fig.10. Given the characteristic mass-splittings of this benchmark, we see that a large portion of the spectrum (consisting of 28 ALPs) in the diphoton distribution could potentially be resolved at the detector. Now, a bin-wise significance estimation, under the conservative assumption of the background events beyond $m_{\gamma\gamma}\approx 80$ GeV, leads to a distribution as shown in Fig.11 with a cumulative significance $S_{c,138}=2.12\sigma$ over the signal region. While spectral techniques Giudice et al. (2018); Beauchesne and Kats (2020) may enhance sensitivity in scenarios where multiple resonances are moderately separated and experimentally resolvable, the present analysis adopts a more conservative, profile-based strategy. This approach focuses on the overall signal shape, which effectively manifests as a single broad structure due to small mass splittings and finite detector resolution.
IV.2.2 Benchmark Point II(b)
@@
For either the intermediate or the heavy-mass cases, the individual peaks are, thus, expected to be visible. The presence of such a bunch of closely packed resonances, with the specific envelope, would immediately point towards a clockwork-like scenario [^27] Naively, an alternative could be a composite sector with multiple excitations. However, not only would such a compositeness scale be very surprising, the abrupt ending of the excitations as well as the shape of the envelope would be hard to explain..
@@
IV.3 Heavy ALPs
For ALPs that are significantly heavier than considered hitherto, the decay photons would be expected to carry sufficient energy (and $p_{T}$) for the event to be triggered even without any additional jet. An example is afforded by the aforementioned Benchmark III, wherein the ALP masses span the range $\sim 150-450$ GeV. However, although extra jets are not needed for triggering, events with such jets do contribiute significantly to the signal cross sections. Hence, we define the signal as a semi-inclusive one composed of a pair of energetic photons with upto two additional jets [^28] The inclusion of events with three or more jets does not improve the signal to noise ratio.. The corresponding $K$-factors, for this mass range, are taken from refs.Ahmed et al. (2016); Boughezal et al. (2015).
@@ IV.3 Heavy ALPs
Table 3: Acceptance cuts for the final state objects in benchmark III Aad et al. (2021).
The background, thus, would receive contributions from essentially the same channels as assumed in the low mass case. The signal profile being different, we, though, would need to adequately tune the assessment, and, in this, we are guided by the ATLAS diphoton analysis Aad et al. (2021) germane to this mass range. Post showering and hadronization, using Pythia8, we take for the identification and isolation of photons in the Delphes detector simulation a cone size of $\Delta R=0.4$ around the photon candidate. The requirement on the parameter $p_{T}^{\rm ratio}$, as defined earlier, is now reset to $p_{T}^{\rm ratio}<0.12$. The complete photon and jet identification criteria, as adapted from the ATLAS analysis ref.Aad et al. (2021), are displayed in Table 3. The $p_{T}$ distributions of the photons and the jets for the generated signal [^29] In contrast with benchmarks I and II, events without jets in the final state also contribute significantly in this case and, therefore, $N(2j)<N(1j)<N_{Tot}$. events are shown in Fig.15 (top), adopting a bin size of $16$ GeV as in the ATLAS searchAad et al. (2021).
Figure 15: BP-III: event distributions, after applying acceptance cuts alone, vide. Table 3. (top) $p_{T}$ distributions for the leading photon ($\gamma_{1}$), sub-leading photon ($\gamma_{2}$), leading jet ($j_{1}$) and the sub-leading jet ($j_{2}$) and (bottom) the two-photon invariant mass.
Fig.15 (bottom) shows the corresponding diphoton invariant mass distribution. Taking cue from ref.Aad et al. (2021), we then apply the selection cuts mentioned in Table 4 in order to define the fiducial volume in the phase space.
@@
$pp\rightarrow a_{n}\,$ $|\eta_{\gamma}|<2.37$ (excluding barrel-to-endcap region $1.37<|\eta_{\gamma}|<1.52$),
@@
Table 4: Event selection cuts applied to form the fiducial signal regions for benchmark III Aad et al. (2021).
The resulting diphoton invariant mass distribution is shown in Fig.16, whereas Fig.17 shows the corresponding bin-wise significance (with the background profile adopted from ref.Aad et al. (2021)) over the diphoton signal region with a cumulative significance $S_{c,138}\approx 1.58\sigma$. For this case, the projected enhancement for a $300\,\rm fb^{-1}$ luminosity reach is $S_{c,300}\approx 2.33\sigma$. Thus, in contrast to benchmarks I and II, the high mass scenario will not be readily accessible in the LHC’s Run 3. It is perhaps worth speculating, though, that benchmark III, with heavy ALPs, could potentially reach the discovery threshold during the forthcoming high luminosity phase of the LHC with a projected luminosity enhancement by a factor of nearly 20 times the current value.
@@
V The Vector-like quarks and the heavy scalars
Beyond the low energy spectrum of the pseudoscalars, the model, as described in section II, also contains a heavy quark with the $SU(3)_{c}\times SU(2)_{L}\times U(1)_{Y}$ charge assignment $(3,1,2/3)$ as well as $N+1$ heavy radial scalar singlets (the partners of the pseudoscalars). Although not germane to the main objectives of our analysis, it is worth outlining here the dynamics of both the VLQ and the heavy scalars so as to establish the consistency of the model, especially in view of the fact that the assumed mass scales (characterised by the SSB scale $f$) of these heavy particles are, in principle, accessible at the LHC.
Considering that $m^{2}\ll f^{2}$ in all the benchmarks, the off-diagonal quadratic terms for the radial scalars in the Lagrangian (eq.(4)) are hierarchically smaller than the diagonal mass terms governed by the coupling $\lambda$, which is presumed to assume $\mathcal{O}(1)$ values. Consequently, all the radial scalars are nearly degenerate with masses $\sim\sqrt{2\lambda}f$. To simplify matters, we assume the other new quartic coupling, viz. $\lambda_{\Phi H}$, is also small, thereby automatically relaxing constraints from the stability of the SM Higgs potential, triviality, etc. Similar to the pseudoscalars, these heavy scalars have effective interactions with the gluons, photons and the EW vector bosons with the respective couplings given by $g_{\pi VV}/\xi$, where $g_{\pi VV}$ represent the pseudoscalar couplings as shown in eq.(17).
@@ V The Vector-like quarks and the heavy scalars
$\begin{split}\left(\Psi_{L}\quad u^{(3)}_{L}\right)^{T}&=U\,\left(T_{L}\quad t_{L}\right)^{T}\ ,\qquad{\rm and}\\ \left(\Psi_{R}\quad u^{(3)}_{R}\right)^{T}&=V\,\left(T_{R}\quad t_{R}\right)^{T}\ ,\end{split}$ (23)
@@
$\begin{split}{\cal L}_{T-V}\displaystyle=\left(\frac{g}{\sqrt{2}}U_{tT}\right)W_{\mu}\,\bar{b}_{L}\gamma^{\mu}T_{L}\;+\left(\frac{g}{2c_{w}}U^{*}_{tt}U_{tT}\right)Z_{\mu}\,\bar{t}_{L}\gamma^{\mu}T_{L}+\displaystyle\frac{g}{c_{w}}\,Z_{\mu}\,\bar{T}\gamma^{\mu}\left[-\,\frac{2}{3}s_{w}^{2}+\frac{g}{2c_{w}}U^{*}_{Tt}U_{tT}P_{L}\right]T+\mbox{h.c.}\end{split}$ (24)
@@
$\begin{split}\mathcal{L}_{T-\phi}&=-\frac{1}{\sqrt{2}}\bar{t}_{L}\left[\left(\lambda_{\Psi}U^{*}_{tT}V_{TT}+y^{\prime}_{\Psi}U^{*}_{tT}V_{tT}\right)\phi_{N}+\left(\lambda_{h}U^{*}_{tt}V_{tT}+y_{\Psi}U^{*}_{tt}V_{TT}\right)h\right]T_{R}\\ &-\frac{1}{\sqrt{2}}\bar{T}_{L}\left[\left(\lambda_{\Psi}U^{*}_{TT}V_{Tt}+y^{\prime}_{\Psi}U^{*}_{TT}V_{tt}\right)\phi_{N}+\left(\lambda_{h}U^{*}_{Tt}V_{tt}+y_{\Psi}U^{*}_{Tt}V_{Tt}\right)h\right]t_{R}+\mbox{h.c.}\ ,\end{split}$ (25)
where $\lambda_{h}$ denotes the top-Higgs Yukawa coupling. To obtain the pseudoscalar FCNCs, it is perhaps the most convenient to make use of the shift symmetry and rescale the field $\Psi_{L}\to e^{i\xi\pi_{N}/f}\Psi_{L}$ in eq.(10). This, in turn, results in a pseudovector FCNC term in the full Lagrangian through the kinetic term of the field $\Psi_{L}$, namely,
$\mathcal{L}_{T-\pi}=\frac{\partial_{\mu}\pi_{N}}{f}\,\bar{T}_{L}\gamma^{\mu}\left(U^{*}_{TT}U_{Tt}\right)t_{L}+\mbox{h.c.},$ (26)
with $\pi_{N}$ being expressed in terms of the mass eigenstates $a_{n}$ through $\pi_{N}=\sum_{k}C^{-1}_{Nk}a_{k}$ with the matrix $C$ as defined in eq.(6). For $f\gtrsim 1000$ GeV and $m_{T}\sim\lambda_{\Psi}f/\sqrt{2}\,>(m_{W,Z,h,\phi}+m_{t})$, the primary decay channels for the VLQ are clearly $T\to b\,W$, $T\to t\,Z$, $T\to t\,h$, $T\to t\,\phi_{N}$ (if kinematically allowed) and $T\to t\,a_{n}$ (see appendix B for the decay width expressions Atre et al. (2011)). Note that the decays to the new scalars and pseudoscalars appear in addition to the conventional VLQ decay modes to the SM particles and, therefore, it is worth comparing the branching fractions pertaining to the different channels. As this needs specifying a few further parameters, we begin by doing so for each of the benchmark points:
BP-I and II: We assume $\lambda_{\Psi}=2.2$, $y_{\Psi}=\epsilon y^{\prime}_{\Psi}$, $\lambda=1.8$ and $y^{\prime}_{\Psi}\lesssim 0.1$ with $\epsilon=0.1$. This choice of $\lambda_{\Psi}$ is motivated by the experimental lower bound on the VLQ mass ($m_{T}\gtrsim 1500$ GeV) to be discussed below. However, as can be expected from such a large value of the coupling, the evolution is fast and an examination of the three-loop renormalization group equations Buttazzo et al. (2013) leads to the conclusion that the theory becomes a very strongly coupled one at a scale $\mu\sim 10$ TeV, signalling that new physics must take over well before this scale [^30] A cutoff near $10$ TeV also serves to mitigate probable issues related to naturalness in the extended scalar sector, exhibiting only a little hierarchy problem.. Given that the clockwork model is not a UV-complete one, this, per se, may still be overlooked. However, we return to this point in the next section.
Note that the aforementioned collider limits assume that the VLQ decays within the detector. If, on the other hand, its decay length was large enough that it does not decay within the LHC detector [^31] This could happen if both $y^{\prime}_{\Psi}$ and $y_{\Psi}$ are tiny. The VLQ, nonetheless, has to decay on cosmological time scales so as not to violate the strong constraints on coloured and electrically charged dark matter candidates. Possibility of a bound state, made of stable neutral VLQs, being a DM candidate is discussed in De Luca et al. (2018)., it would hadronize and leave very distinctive signals. While such “R-hadrons” have been considered in the literature Baer et al. (1999); Arkani-Hamed and Dimopoulos (2005); Giudice and Romanino (2004); Kilian et al. (2005); Hewett et al. (2004); Arvanitaki et al. (2007); Fairbairn et al. (2007); Diaz-Cruz et al. (2007); Choudhury et al. (2008); Di Luzio et al. (2015), the corresponding experimental studies are not as exhaustive ATL (2019). The limits (typically ranging from $\sim 900$ GeV to $\sim 1250$ GeV) depend on the details of the hadronization model assumed. We refrain from considering this possibility in our analysis.
@@
$m_{T}$ lower limit 1540 GeV Tumasyan et al. (2023); Alves et al. (2024) 1500 GeV $\approx$ 1540 GeV
Table 5: VLQ branching fractions for BP-I,$I$ and III (The SM column refers to the assumed branching fractions for VLQ searches Aad et al. (2023c); Tumasyan et al. (2023)). The bottom row displays the resulting lower bounds on $m_{T}$, as derived from the limits on the VLQ pair-production cross-section presented in ref.Tumasyan et al. (2023)
The constraints on the VLQ sector would, thus, come from direct observations at colliders [^32] Collider phenomenology of such KSVZ-VLQs have also been studied in refs.Ghosh et al. (2022); Ghosh and Konar (2023, 2024).. At the large hadron collider, the overwhelmingly leading production mechanism is the QCD-driven one [^33] With $U_{tT}$ not being large, single production is suppressed and the consequent bounds Aad et al. (2023c) are relaxed.. Once pair-produced, the $T$s would decay promptly. Table 5 lists the leading VLQ branching ratios for the three benchmark points [^34] For the chosen set of couplings, a decay of $T$ to radial scalars is kinematically forbidden. For any phenomenologically viable set of couplings, the branching fractions remain much smaller than those listed in the table.. It can be readily ascertained that the branching fractions have a very small dependence on the free parameter $y^{\prime}_{\Psi}$.